OpenStudy (anonymous):

Graph the solution: 2x+3y>-6

OpenStudy (anonymous):

6>2x+3y>-6 3>x+3y/2>-3 3>x+3y/2 x+3y/2>-3 3-3y/2>x x>-3y/2-3 find x and y interceps then calculate y values by putting x for which function is defined and freae a graph can you do it?

OpenStudy (campbell_st):

just graph the line 2x + 3y = - 6 by finding the x intercepts and y intercepts. because of the inequality > the line is dotted...

OpenStudy (anonymous):

Can I substitute x with 0? So like for example. . . .2(0)+3y=-6?

OpenStudy (campbell_st):

thats is y intercept is -2 or y = -2 now let y = 0 so you have 2x + 3(0) = -6

OpenStudy (campbell_st):

can you calculate the x- intercept...?

OpenStudy (anonymous):

-3?

OpenStudy (campbell_st):

great so the line will be dotted and pass through (-3, 0) and (0, -2)  the lins is suppose to be straight... does it make sense...

OpenStudy (campbell_st):

ok... so you needed to shade the right side of the dotted line for the solution.

OpenStudy (anonymous):

How do you know when it's suppose to be dotted?

OpenStudy (campbell_st):

ok... for < or > the line is dotted.. if its greater than or equal to or less than or equal to the line is solid. so the equals really means include any points on the line. to find the side of the line to shade I pick a point, normally (1, 1) and substitute it into the equation so 2(1) + 3(1) >-6 true or false 5 > - 6 is true so (1, 1) is part of the solution  so shade that side of the line.

OpenStudy (anonymous):

Okay i got it!!! Thank you so much! :)

OpenStudy (campbell_st):

glad to help